# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ 4.1 $\mathrm{H}^{1}$-BMO duality
Before we proceed to the second half of Theorem 2, recall that the strong dyadic Christ–Goldberg maximal function corresponding to a weight $W$ on $\mathbb{R}^{n+m}$ (and exponent $p$) is defined as
$M_{\bm{\mathcal{D}},W}f(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|W(x)^{1/p}f|\rangle_{R}\mathbf{1}_{R}(x),\qquad x\in\mathbb{R}^{n+m},\leavevmode\nobreak\ f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n+m};\mathbb{C}^{d}).$
Because of [Vuorinen_2023, Theorem 1.3] we have that the operator $M_{\bm{\mathcal{D}},W}$ is bounded when acting on $L^{p}(W)\rightarrow L^{p}(\mathbb{R}^{n+m})$ (note that the target space is unweighted) provided that $[W]_{A_{p},\bm{\mathcal{D}}}<\infty,$ specifically one has the bound
$\|M_{\bm{\mathcal{D}},W}\|_{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p}}^{2/(p-1)},$ (4.2)
for $1<p<\infty.$
@@
$\ell(\Phi)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\forall\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$
@@
We partially adapt the second half of the proof of [Isralowitz_2017, Theorem 1.3]. We denote by $\mathcal{H}$ the set of all $\mathrm{M}_{d}(\mathbb{C})$-valued $\mathrm{L}^{2}$ functions on $\mathbb{R}^{n+m}$ with finitely many nonzero biparameter cancellative Haar coefficients, and we identify such functions with finitely supported sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ in the obvious way. Then, for all $\Phi\in\mathcal{H},$ we have
@@
$\ell(\Phi)$ $=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)E_{ij})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)\ell(E_{ij})$
$=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})=(\Phi,B),$
where $E_{ij}$ is the $d\times d$-matrix with 1 at the $(i,j)$-entry and 0 at all other entries, $A(i,j)$ is the $(i,j)$-entry of a matrix $A,$ and
$B_{R}^{\varepsilon}(i,j)\coloneq\overline{\ell(h_{R}^{\varepsilon}E_{ji})},\qquad\forall i,j=1,\ldots,d,\leavevmode\nobreak\ \forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall\varepsilon\in\mathcal{E}.$
By Proposition 8 and since $\mathcal{H}$ is dense in $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),$ it suffices only to prove that the defined sequence $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ is in $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)$ with
@@
$\sim_{n,m,p,d}\sup_{\Omega}\frac{1}{|\Omega|^{1/p^{\prime}}}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m})}=:\mathcal{C},$
@@
[…]
@@
$\sim_{n,m,p,d}\left\{\left|\int_{\mathbb{R}^{n+m}}\langle F(x),G(x)\rangle\,\mathrm{dm}(x)\right|:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\right\}$
=$=\bigg{\{}\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((G^{\varepsilon}_{R})^{\ast}F_{R}^{\varepsilon})\bigg{|}:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\bigg{\}},$ (4.4)
@@
[…]
@@
(4.1)${\sim}_{n,m,p,d}\bigg{\|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}h_{R}^{\varepsilon}\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
@@
$=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|(\hat{A},B)|=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|\ell(\hat{A})|$
@@
$\hat{A}^{\varepsilon}_{R}\coloneq\begin{cases}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1},\text{ if }R\in\bm{\mathcal{D}}(\Omega)\\ \\ 0,\text{ otherwise}\end{cases},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
@@
We define now functions $\widetilde{N}_{\Omega}^{(k)}$ by applying the strong dyadic Christ–Goldberg maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$). That is, we define
@@
because the scalar weight $|V^{1/p}M|^{p}$ is uniformly in the biparameter $\bm{\mathcal{D}}$-dyadic Muckenhoupt $A_{p}$ class for every positive definite matrix $M$ (see for example [DKPS2024, Lemma 3.4]). Next, using the comparability between the matrix norm and the supremum of norms of matrix columns, we get
$N_{\Omega}(x)$ $\lesssim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\sup_{k=1,\ldots,d}\langle|V(x)^{-1/p}V^{1/p}e_{k}|\rangle_{R}\mathbf{1}_{R}(x)$
$\leq[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\widetilde{N}_{\Omega}(x).$
Finally, using the boundedness (4.2) of the strong dyadic Christ–Goldberg maximal operator, we obtain
$\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}.$
@@
$\lesssim_{n,m,p,d}\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
@@
It is clear that the proofs of Propositions 8, 9 work also in the one-parameter setting.
@@ 4.2 Two-matrix weighted bounds for paraproducts
Let $B:\mathbb{R}^{n+m}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|\bigg{)}^{1/2},$
where $\{B_{R}^{\varepsilon}\}$ is the sequence of biparameter Haar coefficients of $B$. Following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “pure” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
@@
$\Gamma_{\bm{\mathcal{D}},B}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{1\oplus\varepsilon\oplus\delta}B_{R}^{\varepsilon}f_{R}^{\delta}.$
@@
and extend these operations in the obvious entrywise fashion to $\mathcal{E}^{1},$ $\mathcal{E}^{2}$ and $\mathcal{E}.$ Moreover, abusing notation we denote $(1,\ldots,1)$ (where the number of entries is always clear from the context) by just $1$ Notice that $1\oplus\varepsilon_{i}\oplus\delta_{i}\neq 1,$ for all $\varepsilon,\delta\in\mathcal{E}$ with $\varepsilon_{i}\neq\delta_{i},$ $i=1,2.$
@@
$\Gamma_{\bm{\mathcal{D}},B}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{\varepsilon}B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\delta},$
@@
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
@@
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
@@
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,p,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt part of the proof of [Isralowitz_2017, Theorem 2.2]. Note that by the John–Nirenberg inequalities for (unweighted) scalar dyadic product BMO we have
@@
Let us first see that $\mathcal{C}\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$ Let $\Omega$ be any Lebesgue-measurable subset of $\mathbb{R}^{2}$ of nonzero finite measure. Let also $e\in\mathbb{C}^{d}\setminus\{0\}$ be arbitrary. We test $\Pi_{\bm{\mathcal{D}},B}^{(11)}$ on the function $f\coloneq\mathbf{1}_{\Omega}U^{-1/p}e.$ Using Lemma 5.3 of [DKPS2024] we obtain
@@
$\geq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}.$
Note also that $\|f\|_{\mathrm{L}^{p}(U)}=|\Omega|^{1/p}\cdot|e|.$ Therefore, we see that
$\frac{1}{|\Omega|^{1/p}}\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}|e|,$
for all $e\in\mathbb{C}^{d}.$ In view of (2.3) we deduce
@@
By part (2) of Lemma 3 we have
$|P\langle U^{-1/p}\rangle_{R}|=|\langle U^{-1/p}\rangle_{R}P^{\ast}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P|\sim_{p,d}|\mathcal{U}^{-1}_{R}P|=|P\mathcal{U}^{-1}_{R}|,\quad\forall P\in\mathrm{M}_{d}(\mathbb{C}),$

### v3

@@ 4.1 $\mathrm{H}^{1}$-BMO duality
Before we proceed to the second half of Theorem 2, recall that the dyadic biparameter Christ–Goldberg maximal operator corresponding to a weight $W$ on $\mathbb{R}^{n+m}$ (and exponent $p$) is defined as
$M_{\bm{\mathcal{D}},W}f(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|W(x)^{1/p}f|\rangle_{R}\mathbf{1}_{R}(x),\qquad x\in\mathbb{R}^{n+m},\penalty 10000\ f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n+m};\mathbb{C}^{d}),$
and the dyadic biparameter Christ–Goldberg auxiliary maximal operator as
$\widetilde{M}_{\bm{\mathcal{D}},W}f\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|\mathcal{W}_{R}f|\rangle_{R}\mathbf{1}_{R},\qquad x\in\mathbb{R}^{n+m},\penalty 10000\ f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n+m};\mathbb{C}^{d}).$
Because of [41, Theorem 1.3] we have that the operator $M_{\bm{\mathcal{D}},W}$ is bounded when acting on $\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})$ (note that the target space is unweighted) provided that $W]_{A_{p},\bm{\mathcal{D}}}<\infty,$ specifically one has the bound
$\|M_{\bm{\mathcal{D}},W}\|_{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{2/(p-1)},$ (4.8)
for $1<p<\infty.$ From [14, Proposition 4.1] we also have
$\|\widetilde{M}_{\bm{\mathcal{D}},U}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{(p+1)/(p(p-1))}.$
@@
$\ell(\Phi)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\forall\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$ (4.9)
@@
We partially adapt the second half of the proof of [28, Theorem 1.3]. The main difference is that we need to introduce an auxiliary function $N_{\Omega}$ and the properties of scalar weights of the form $|V^{1/p}A|^{p}$ with positive definite matrices $A$ in order to get the necessary bounds. There is also an extra twist given in (4.1) below.
We denote by $\mathcal{H}$ the set of all $\mathrm{M}_{d}(\mathbb{C})$-valued $\mathrm{L}^{2}$ functions on $\mathbb{R}^{n+m}$ with finitely many nonzero biparameter cancellative Haar coefficients, and we identify such functions with finitely supported sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ in the obvious way. In addition, given a matrix $A$ we denote here its $(i,j)$ entry by $A(i,j).$ Then, for all $\Phi\in\mathcal{H},$ we have
@@
$\ell(\Phi)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)E_{ij})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)\ell(E_{ij})$ (4.10)
where $E_{ij}$ is the $d\times d$-matrix with 1 at the $(i,j)$-entry and 0 at all other entries. Thus, if we take $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ given by
$B_{R}^{\varepsilon}(i,j)\coloneq\overline{\ell(h_{R}^{\varepsilon}E_{ji})},\qquad\forall i,j=1,\ldots,d,\penalty 10000\ \forall R\in\bm{\mathcal{D}},\penalty 10000\ \forall\varepsilon\in\mathcal{E},$
the last expression of (4.10) is exactly (4.9). By Proposition 7 and since $\mathcal{H}$ is dense in $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),$ it suffices only to prove that sequence $B$ is in $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)$ with
@@
$\sim_{n,m,p,d}\sup_{\Omega}\frac{1}{|\Omega|^{1/p^{\prime}}}\bigg\|\bigg(\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg)^{1/2}\bigg\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m})}=:\mathcal{C},$ (4.12)
@@
Let us denote by $\langle\,\cdot\,$$\cdot\,\rangle$ the Hermitian product on $\mathrm{M}_{d}(\mathbb{C})$ given by
@@
$\sim_{n,m,p,d}\left\{\left|\int_{\mathbb{R}^{n+m}}\langle F(x),G(x)\rangle\,\mathrm{dm}(x)\right|:\penalty 10000\ G\in\mathcal{H},\penalty 10000\ \|G\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\right\}$
$=\bigg\{\bigg|\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((G^{\varepsilon}_{R})^{\ast}F_{R}^{\varepsilon})\bigg|:\penalty 10000\ G\in\mathcal{H},\penalty 10000\ \|G\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\bigg\},$ (4.13)
@@
$\sim_{n,m,p,d}\bigg(\int_{\mathbb{R}^{n+m}}\bigg(\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|F_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg)^{p/2}dx\bigg)^{1/p},\qquad\forall F\in\mathcal{H}.$ (4.14)
@@
[…]
@@
[…]
@@
$\hat{A}^{\varepsilon}_{R}\coloneq\begin{cases}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1},\text{ if }R\in\bm{\mathcal{D}}(\Omega)\\ \\ 0,\text{ otherwise}\end{cases},\quad R\in\bm{\mathcal{D}},\penalty 10000\ \varepsilon\in\mathcal{E}.$
@@
We define now functions $\widetilde{N}_{\Omega}^{(k)}$ by applying the dyadic Christ–Goldberg biparameter maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$). That is, we define
@@
because the scalar weight $|V^{1/p}M|^{p}$ is uniformly in the biparameter $\bm{\mathcal{D}}$-dyadic Muckenhoupt $A_{p}$ class for every positive definite matrix $M$ (see for example [14, Lemma 3.4]). Next, using the comparability between the matrix norm and the supremum of norms of matrix columns, we get that the previous expression is bounded above by
$V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\sup_{k=1,\ldots,d}\langle|V(x)^{-1/p}V^{1/p}e_{k}|\rangle_{R}\mathbf{1}_{R}(x)\leq[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\widetilde{N}_{\Omega}(x).$
Finally, using the boundedness (4.8) of the dyadic Christ–Goldberg biparameter maximal operator, we obtain
$\|N_{\Omega}\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}.$
@@
$\lesssim_{n,m,p,d}\|N_{\Omega}\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{2})}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
@@
The proofs of Propositions 7, 8 work also in the one-parameter setting.
@@ 4.2 Two-matrix weighted bounds for paraproducts
For a given locally integrable function $B:\mathbb{R}^{n+m}\to\mathrm{M}_{d}(\mathbb{C}),$ we define its $\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}$ norm as the norm of the sequence $\{B_{R}^{\varepsilon}\}$ of its biparameter Haar coefficients. Following the terminology of Holmes–Petermichl–Wick [23, Subsection 6.1] we define the following so-called “pure” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
@@
$\Gamma_{\bm{\mathcal{D}},B}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{1\oplus\varepsilon\oplus\delta}B_{R}^{\varepsilon}f_{R}^{\delta}.$
@@
and extend these operations in the obvious entrywise fashion to $\mathcal{E}^{1},$ $\mathcal{E}^{2}$ and $\mathcal{E}.$ Moreover, abusing notation we denote $(1,\ldots,1)$ (where the number of entries is always clear from the context) by just $1.$ Notice that $1\oplus\varepsilon_{i}\oplus\delta_{i}\neq 1,$ for all $\varepsilon,\delta\in\mathcal{E}$ with $\varepsilon_{i}\neq\delta_{i},$ $i=1,2.$
@@
$\Gamma_{\bm{\mathcal{D}},B}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{\varepsilon}B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\delta},$
@@
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}}.$
@@
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}}.$
@@
Throughout the proof $\lesssim,\gtrsim,\sim$ mean that all implied inequality constants depend only on $n,m,p,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $V]_{A_{p},\bm{\mathcal{D}}},$ unless otherwise specified.
We adapt part of the proof of [28, Theorem 2.2]. Note that by the John–Nirenberg inequalities for (unweighted) scalar dyadic product BMO we have
@@
Let us first see that $\mathcal{C}\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$ Let $\Omega$ be any Lebesgue-measurable subset of $\mathbb{R}^{2}$ of nonzero finite measure. Let also $e\in\mathbb{C}^{d}\setminus\{0\}$ be arbitrary. We test $\Pi_{\bm{\mathcal{D}},B}^{(11)}$ on the function $f\coloneq\mathbf{1}_{\Omega}U^{-1/p}e.$ Using [14, Lemma 5.3] we obtain
@@
$\geq\bigg(\int_{\mathbb{R}^{n+m}}\bigg(\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg)^{p/2}\,\mathrm{dm}(x)\bigg)^{1/p}.$
Note also that $\|f\|_{\mathrm{L}^{p}(U)}=|\Omega|^{1/p}\cdot|e|.$ Therefore, the previous quantity is bounded above by $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}|\Omega|^{1/p}|e|$ for all $e\in\mathbb{C}^{d}.$ Using repeatedly the elementary estimates (3.1) and (3.3) in [14] we conclude
@@
To continue the computation, we need the following fact proved in [25, Lemma 2.2]: If $\mathcal{W}_{E}$ is the reducing operator of a matrix weight $W$ over a set $E$ with respect to exponent $p,$ and $W^{-1/(p-1)}$ is also integrable over $E$ with reducing exponent $\mathcal{W}^{\prime}_{E}$ with respect to exponent $p^{\prime}$, then
$|{\langle W^{1/p}\rangle}_{E}e|\leq|\mathcal{W}_{E}e|\lesssim_{p,d}|\mathcal{W}_{E}\mathcal{W}_{E}^{\prime}|^{d}{\langle W^{1/p}\rangle}_{E}e|,\quad\forall e\in\mathbb{C}^{d}.$ (4.15)
Using (4.15) we have
$|P\langle U^{-1/p}\rangle_{R}|=|\langle U^{-1/p}\rangle_{R}P^{\ast}|\sim|\mathcal{U}^{\prime}_{R}P^{*}|\sim|\mathcal{U}^{-1}_{R}P^{*}|=|P\mathcal{U}^{-1}_{R}|,\quad\forall P\in\mathrm{M}_{d}(\mathbb{C}),$